The Cecil-Booker Vending Company changed its method of valuing inventory from the average cost method to the FIFO cost method at the beginning of 2021. At December 31, 2020, inventories were $115,000 (average cost basis) and were $119,000 a year earlier. Cecil-Booker’s accountants determined that the inventories would have totaled $145,000 at December 31, 2020, and $150,000 at December 31, 2019, if determined on a FIFO basis. A tax rate of 25% is in effect for all years.

One hundred thousand common shares were outstanding each year. Income from continuing operations was $350,000 in 2020 and $475,000 in 2021. There were no discontinued operations either year.

Required:
1. Prepare the journal entry at January 1, 2021, to record the change in accounting principle. (All tax effects should be reflected in the deferred tax liability account.)
2. Prepare the 2021–2020 comparative income statements beginning with income from continuing operations (adjusted for any revisions). Include per share amounts.

Answer is not complete.

Complete this question by entering your answers in the tabs below.

  • Required 1
  • Required 2

Prepare the journal entry at January 1, 2021, to record the change in accounting principle. (All tax effects should be reflected in the deferred tax liability account.) (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)

No Date General Journal Debit Credit
1 January 01, 2021 Inventoryselected answer correct 30,000selected answer correct not attempted
Income tax payableselected answer correct not attempted 13,600 incorrect
Retained earningsselected answer correct not attempted 20,400 incorrect
  • Required 1
  • Required 2

Prepare the 2021–2020 comparative income statements beginning with income from continuing operations (adjusted for any revisions). Include per share amounts. (Round EPS answers to 2 decimal places.)

COMPARATIVE INCOME STATEMENTS
2021 2020
not attempted $475,000selected answer correct $349,000selected answer correct
not attempted 190,000selected answer incorrect 139,600selected answer incorrect
not attempted $285,000 $209,400
Earnings per common share $2.33selected answer incorrect $2.09selected answer incorrect
Answer #1

Journal Entries

Particulars Debit Credit
Inventory (150000-119000) 31000
To Income tax payable (31000*25%) 7750
To Retained earnings 23250

Comparative Income Statement

Particular 2021 2020
Income 475000 350000
Less Income tax (118750) (87500)
(475000*25%) (350000*25%)
net income 356250 262500
EPS =356250/100000=3.56 =262500/100000=2.62